1. The Cartesian coordinates representation of a vector is (65 cm/s, 32 cm/s), the polar coordinates...

1.

The Cartesian coordinates representation of a vector is (65 cm/s, 32 cm/s), the polar coordinates of this vector are:

Select one:

a. (72 cm/s, 26 degrees)

b. (52 cm/s, 63 degrees)

c. (45 cm/s, 15 degrees)

d. (26 cm/s, 72 degrees)

2. Which of the following correctly expresses the vector 29 m at 29 degrees in unit vector notation

Select one:

a. (15 i + 43 j) m

b. (25 i + 14 j) m

c. (14 i + 25 j) m

d. (29 i + 29 j) m
Answer #1

- Corresponding polar con ordinates of Ve Vy) is (Vo) Pe (Vx, vy) Where, V =√x+vy ok a → and o=tan(Y/N) here, Vx = 65 cm/s, V